1. A drag racer constantly accelerates at a rate of 39 m/s2

to increase its speed in a short amount of time. If the drag racer started at rest, what distance did it have to travel to reach a velocity of 33 m/s, E?




1
Expert's answer
2021-10-14T11:31:28-0400
"v^2=v_0^2+2ad,""d=\dfrac{v^2}{2a}=\dfrac{(33\ \dfrac{m}{s})^2}{2\cdot39\ \dfrac{m}{s^2}}=14\ m."